On January 2016, Profit Ltd enters the contract with City Ltd for the use of a truck. The contract contained the following clauses:

Lease term: four years with a $6,000 purchase option at the end of the lease.

Lease payments: $200,048 per annum in advance, the first payment is made on 1 January 2016.

Fair value of the truck : $720,000

Profit Ltd incurs $50,000 initial direct cost, including $15,000 legal fees and $35,000 commission to real estate agent

At the end of the lease, Profit Ltd has express intention to exercise the purchase option.

Implicit interest rate: 8%

Estimated useful life of the truck is five years

Profit Ltd accounts for its lease using IFRS 16

(a) Determine the lease liability and initial measurement of a right-use-use asset (i.e. the truck) for Profit ltd.

(b) Prepare the lease liability movement schedule for each year for Profit Ltd.

(c) Prepare accounting journal entries for Profit Ltd for the years ended 31 December 2016 and 2017
Answer #1

1. Determination of lease liability and initial measurement of leae asset

Payment date Amount Discounting factor @8% Discounted Value
1.1.2016 200,048 1 200,048
1.1.2017 200,048 0.9259 185,224.44
1.1.2018 200,048 0.8573 171,501.15
1.1.2019 200,048 0.7938 158,798.10
31.12.2019 6,000 0.7350 4,410
Lease liability to be recognised 719,981.69

Right to use asset = Lease liability + initial direct costs

= 719,981.69+50,000= 769,981.69

2. Lease liability movement schedule

Date Lease liability (1) Interest @ 8% (2) Payment made (3) Interest Payment (4) Principal Payment (5=3-4) Closing lease liability as on date (6=1-3)
1.1.2016 719,981.69 0 200,048 200,048 519,933.69
31.12.2016 519,933.69 41,594.70 0 0 0 561,528.39
1.1.2017 561,528.39 0 200,048 41,594.7 158,453.3 361,480.39
31.12.2017 361,480.39 28,918.43 0 0 0 390,398.82
1.1.2018 390,398.82 0 200,048 28918.43 171,129.57 190,350.82
31.12.2018 190,350.82 15,228.07 0 0 0 205,578.89
1.1.2019 205,578.89 0 200,048 152,28.07 184,819.93 5,530.89
31.12.2019 5,530.89 469.11 (bal fig) 6,000 469.11 5,530.89 0

3. Journal entries

Date Particulars Debit Credit
1.1.2016 Right to use asset Dr 719,981.69
To lease liability 719,981.69
(lease liability and lease asset recorded)
1.1.2016 Right to use asset Dr 50,000
To Cash 50,000
(Payment of initial direct cost)
1.1.2016 Lease Liability Dr 200,048
To Cash/ Bank 200,048
(lease payment made)
31.12.2016 Interest Expense Dr 41,594.7
To lease liabilty 41,594.7
(Recording of interest)
31.12.2016 Depreciation Dr 192,495.42
To right to use asset 192,495.42
(Depreciation of right to use asset)
1.1.2017 Lease Liabilty Dr 200,048
To Cash/ Bank 200,048
(Payment of lease liability)
31.12.2017 Interest Expense Dr 28,918.43
To cash / bank 28,918.43
(Recording of interest)
31.12.2017 Depreciation Dr 192,495.42
To right to use asset 192,495.42
(Depreciation of right to use asset)

Depreciation will be charged over lesser of lease term or expected useful life i.e. 4 years

Amount of depreciation = 769,981.69/4= 192495.42

31.12.2016 Lease Liability Dr 158,453,3
Interest Expense Dr 41,594.7
To cash / bank 200,048
(Payment of interst and lease liability)
31.12.2016 Depreciation Dr 192,495.42
To right to use asset 192,495.42
(Depreciation of right to use asset)